OpenStudy (anonymous):

Please help me solve for r! r^2-4r+8=5r

OpenStudy (anonymous):

First you should subtract 8. In order to pass it to the other side of the equation.

OpenStudy (anonymous):

Then it should look like this r^2-4r=5r-8 Then subtract the 5r. r^2-4r-5r=-8 Once you did that you can combine the r's. so... r^2-9r=-8

OpenStudy (anonymous):

I did that and I got r^2 -9r +8=0 What I have so far is: 

OpenStudy (anonymous):

Okay, so next you should try to simplify the sqrt.

OpenStudy (anonymous):

I thought 113 was prime

OpenStudy (anonymous):

Then if it can't be simplified any more, I'm pretty sure that's your answer. I know It looks big.

OpenStudy (anonymous):

But the actual answers are 8 and 1

OpenStudy (anonymous):

Like the book tells me the answers are 8 and 1 and r^2 - 9 + 8 = 0 factors out to (r-8)(r-1)

OpenStudy (anonymous):

Hmm... then I'm not too sure sorry.

OpenStudy (anonymous):

Hold on a sec

OpenStudy (anonymous):

if i subtracted r^2 - 4r +8 onto the side 5r is on, would i get a different answer?

OpenStudy (loser66):

minus both sides by 5r what do you have

OpenStudy (loser66):